OpenStudy (anonymous):

solve the following equation (5x - 1)/2 + (-3x - 1)/4 = 9/2 any help works

OpenStudy (maheshmeghwal9):

take common denominator "4" & just do add & solve the eq.

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

\[\frac{5x-1}{2}+\frac{-3x-1}{4}=\frac{9}{2}\] Multiply each part by a form of 1 (written as (1/1), etc.

OpenStudy (anonymous):

\[\frac{5x - 1}{2} + \frac{-3x - 1}{4} = \frac{9}{2}\]Common denominator is 4, so get the common denominator by multiplying by values of 1.\[\frac{5x - 1}{2} \times \frac22+ \frac{-3x - 1}{4} = \frac{9}{2} \times \frac22\]\[\frac{2(5x - 1)}{4} + \frac{-3x - 1}{4} = \frac{18}{4}\]\[10x - 2 - 3x - 1 = 18\]I trust that you can combine like terms and solve for x now that the denominator is gone?

OpenStudy (anonymous):

Once again thank you.

OpenStudy (anonymous):

No Problem :)

OpenStudy (anonymous):

@golikrishna please don't give out the answers